OpenStudy (anonymous):

Solve for y: 3x – 4y = 8

OpenStudy (anonymous):

y= (3x-8)/4

OpenStudy (anonymous):

get y on its own

OpenStudy (anonymous):

3x -4y = 8 3x =8 + 4y 3x - 8 =4y y= (3x -8)/4